OpenStudy (anonymous):

i will always fan and medal when given help The mean weight of the adult population of a country is 67.4 kg, with a standard deviation of 8 kg. If a person is chosen at random, find the probability that that person weighs more than 85 kg.

OpenStudy (kirbykirby):

So let \(X\) be a person's weight. The problem asks you to find \(P(X>85)\).Assuming X is normally distributed, you can standardize X: \[Z=\frac{X-\mu}{\sigma} \] where \(\mu\) is the mean and \(\sigma\) is the standard deviation. \[P(X > 85) = P\left( \frac{X-67.4}{8}>\frac{85-67.4}{8}\right)=P(Z>2.2)\]

OpenStudy (anonymous):

thanksss so much

OpenStudy (kirbykirby):

:)